40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 10 19:59:09 2020, Yehonathan, Update, Loss Measurement, Arm transfer function measurement YArmFrequencyResponse.pdf
    Reply  Thu Mar 12 10:43:50 2020, rana, Update, Loss Measurement, Arm transfer function measurement 
       Reply  Mon Mar 16 15:23:03 2020, Yehonathan, Update, Loss Measurement, Arm transfer function measurement Cross_calibration_setup.jpgPDA520overPDA255_Response.pdfYArmFrequencyResponse.pdf
          Reply  Sat Apr 18 14:57:44 2020, Yehonathan, Update, Loss Measurement, Arm transfer function measurement 
Message ID: 15264     Entry time: Tue Mar 10 19:59:09 2020     Reply to this: 15269
Author: Yehonathan 
Type: Update 
Category: Loss Measurement 
Subject: Arm transfer function measurement 

I want to measure the transfer function of the arm cavities to extract the pole frequencies and get more insight into what is going on with the DC loss measurements.

The idea is to modulate the light using the PSL AOM. Measure the light transmitted from the arm cavities and use the light transferred from the IMC as a reference.

I tried to start measuring the X arm but the transmission PD is connected to the QPD whitening filter board with a 4 pin Lemo for which I couldn't find an adapter.

  • I switch to the Y arm where the transmission PD - Thorlabs PDA520 (250KHz Bandwidth) - is BNC all the way.
  • I lay an 82ft BNC cable from the Y Arm 1Y4 to 1Y1 where the BNC from the IMC Trans PD and an SR785 are found. 
  • I lock the Arm cavities.
  • I connect the AOM cable to the source, the TRY PD (Teed off from the QPD whitening filter) to CH1 and IMC_TRANS to CH2 and measure the transfer function using a swept sine with an offset of 300mV and amplitude of 100mV.
  • I fit it to a low pass filter function - see attachment 1 - but it seems like the fit rails off at 10KHz. 

Could this be because of the PDA520 limited BWs? I tried playing with the PD gain/bandwidth switch but it seems like it was already set to high bandwidth/low gain.

In any case, the extracted pole frequency ~ 2.9kHz implies a finesse > 600 (assuming FSR = 3.9MHz) which is way above the maximal finesse (~ 450) for the arm cavities.

I disconnected the source from the AOM. But left the other two BNCs connected to the SR785. Also, TRY PD is still teed off. Long BNC cable is still on the ground.

Attachment 1: YArmFrequencyResponse.pdf  26 kB  Uploaded Tue Mar 10 22:16:45 2020  | Hide | Hide all
YArmFrequencyResponse.pdf
ELOG V3.1.3-